OpenStudy (shelby1290):

At his summer job, Sam's regular wage is $9.50/h. For any overtime, Sam earns 1.5 times his regular wage. a) write an equation representing Sam's regular pay b) write a seperate equation representing sam's overtime pay c) sam gets a raise to $10/h. how does this change affect the equations?

OpenStudy (anonymous):

For a, we do not care about the overtime His wage is 9.50 for every hour, so lets call this function f(h). A function of the hours he did. f(h)=9.50 * h

OpenStudy (anonymous):

Can you try to do the rest? It is very similar logic

OpenStudy (shelby1290):

@Andras i dont think i can

OpenStudy (shelby1290):

but do you need the overtime part for part b

OpenStudy (shelby1290):

or do i just use the same equation or something

OpenStudy (shelby1290):

i dont really get the rest of it

OpenStudy (shelby1290):

^ the question

OpenStudy (anonymous):

The overtime is 1.5 times: 9.5*1.5=14.25 So lets call that function g(h) g(h)=14.25h

OpenStudy (shelby1290):

okay

OpenStudy (anonymous):

sam gets a raise to $10/h. how does this change affect the equations? Same as before just replace 9.5 with 10

OpenStudy (shelby1290):

so then would i just write g=10h

OpenStudy (shelby1290):

or 10x1.5

OpenStudy (anonymous):

f(h)=9.50 * h changes to? 9.5*1.5=14.25 g(h)=14.25h changes to?

OpenStudy (shelby1290):

ohhhhh so substitue 10 for all of the equations

OpenStudy (shelby1290):

right?

OpenStudy (anonymous):

yep, so what is your new f(h) and g(h)?

OpenStudy (shelby1290):

f(h)=10xh

OpenStudy (anonymous):

and?

OpenStudy (shelby1290):

10x1.5=15

OpenStudy (shelby1290):

g(h)=15h

OpenStudy (anonymous):

yes, you got that right :)

OpenStudy (shelby1290):

so do i need to explain in words why it will affect it or do i just write that the outcome will be different

OpenStudy (shelby1290):

..or leave the new equations there

OpenStudy (shelby1290):

with their functions?

OpenStudy (anonymous):

write down the new equations, that should do. maybe label the new functions a bit differently Lets say instead of f(h) and g(h) you can write f*(h) and g*(h)

OpenStudy (shelby1290):

oh okay and thanks for the help!

OpenStudy (anonymous):

Welcome